Solve the system by graphing. (If the system is inconsistent, enter INCONSISTENT. If the system is dependent, enter DEPENDENT.)
x - y = -1
2x + 3y = 18
Solve the system by graphing. (If the system is inconsistent, enter INCONSISTENT. If the system is dependent, enter DEPENDENT.)
4x + 2y = 18
5x - 4y = 3

Answer by TimothyLamb(4379)

You can graph these yourself. I will give you the correct answers:
---
x - y = -1
2x + 3y = 18
---
put the system of linear equations into standard form
---
x - y = -1
2x + 3y = 18
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 3
y = 4
---
4x + 2y = 18
5x - 4y = 3
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 3
y = 3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---